Hello, Could you help me on the solution to this question,
The sum of a series is given by the formula Sn=(4^n)(n-3). Determine the value of the 6th term.
Thanks all!.

Answer by stanbon(75887)

The sum of a series is given by the formula Sn=(4^n)(n-3). Determine the value of the 6th term.
-------
The 6th term = the sum of 6 terms - the sum of 5 terms.
6th term = (4^6)(6-3) - 4^5(5-3)
=4^5(4(3)-(5-3))
=1024(10)
=10240
Cheers,
Stan H.